OpenStudy (anonymous):

integrate from -infinity to infinity. xe^((-x)^2)

OpenStudy (anonymous):

i know that it turns out to be \[\frac{ -1 }{ 2 } e^{-x^{2}}\]

OpenStudy (anonymous):

and im not sure if i need to split up the boundaries to -infinity to 0 and then 0 to infinity.

OpenStudy (freckles):

yep you need a splitting

OpenStudy (anonymous):

i know when i split, the equation can't be improper, by splitting on 0 is that okay?

OpenStudy (freckles):

you can choose to split it up at 0 or at any constant number between -inf and inf

OpenStudy (anonymous):

and , when i try to evaulate it from -infintiy to 0 i get -infinity, and if i evalulate it from 0 to infinity i get +infinity

OpenStudy (anonymous):

would that be 0 or would i have to use l'hopital?

OpenStudy (anonymous):

or am i wrong in my evaluation.

OpenStudy (freckles):

We need to evaluate this limit \[\frac{-1}{2}e^{-x^2}|_{-\infty}^{0}\] and also \[\frac{-1}{2}e^{-x^2}|_{0}^{\infty}\] If both limits converge (say one to a and the other to b) then the sum of these limits will converge to a+b If one of diverges, then the sum diverges.

OpenStudy (freckles):

I don't even need l'hospital for this x^2 is positive -x^2 is negative e^(really negative large numbers) approaches 0

OpenStudy (freckles):

but don't forget about the constant parts up there ...

OpenStudy (anonymous):

the first evaluation is -infinity because its -1/2 -infinity, and the second one is 0 +1/2

OpenStudy (freckles):

They should both converge

OpenStudy (freckles):

first one is -1/2-0 second one is 0+1/2

OpenStudy (anonymous):

is it because you're squaring the -infinity first before accounting the -x

OpenStudy (freckles):

squaring negative numbers is still positive

OpenStudy (anonymous):

so like -(-infinity^2)

OpenStudy (freckles):

you mean -(-infinity)^2

OpenStudy (anonymous):

i guess its order of operation LOL

OpenStudy (freckles):

Pretend f(x)=x^2 f(-x)=f(x)

OpenStudy (freckles):

f is an even function

OpenStudy (freckles):

Like pluggin in -1 and 1 both give the same f which is 1

OpenStudy (freckles):

\[g(x)=-x^2 \]

OpenStudy (freckles):

This is still an even function

OpenStudy (freckles):

pluggin in -1 or 1 still gives you -1

OpenStudy (anonymous):

i understand now :) thank you for your help. im studying for an exam so if you wanna stick around i'll have more questions

OpenStudy (freckles):

It is actually really late here...But I might be on for like a few more minutes